OpenStudy (anonymous):

If sin(theta) =4/5 and is in quadrant 2, the value of cot(theta)....Help Dx

OpenStudy (anonymous):

In order to find cot(theta) you need to know \[\cot(\theta)\] in addition to \[\sin(\theta)\].. correct? Since your \[\cot(\theta) = \cos(\theta)/\sin(\theta)\] ....

OpenStudy (anonymous):

so far you have \[\sin(\theta)=4/5\], and you can use that to find that \[\cos(\theta)=3/5\] Once you have that you can just plug in to find \[\cot(\theta)\]

OpenStudy (anonymous):

so what would \[\cot(\theta) = \] ?

OpenStudy (anonymous):

mistake on my first post.. I meant to say "cos in addition to sin" ... but does this all make sense to you?

OpenStudy (anonymous):

-3/4 ?

OpenStudy (anonymous):

yeah, actually you would have a (-3/5) for the cos .. you're right.. and then you would get -3/4 for cot

OpenStudy (anonymous):

since we're in the second quadrant

OpenStudy (anonymous):

Ok thanks :)